OpenStudy (anonymous):

Use sec x= -3/2 and tan x>0 to find the values of all six trigonometric values

OpenStudy (anonymous):

well (tanx)^2+1=(secx)^2 use it to find tanx also cosx=1/secx as tanx >0 so its in 3rd quadrant where cosx, secx. cosecx. sinx are negative and cotx, tanx are positive sinx=(1-(cosx)^2)^(1/2) also cosecx =1/sinx and cotx=1/tanx use these to find rest 6 trigonometric functions

OpenStudy (anonymous):

Thank you!:)